The angular position of a point on a rotating wheel is given by θ = 3.04 + 4.83t2 + 2.99t3, where θ is in radians and t is in seconds. At t = 0, what are (a) the point's angular position and (b) its angular velocity? (c) What is its angular velocity at t = 7.43 s? (d) Calculate its angular acceleration at t = 1.12 s. (e) Is its angular acceleration constant?


1
Expert's answer
2022-03-20T18:50:55-0400

Given:

"\theta = 3.04 + 4.83t^2 + 2.99t^3"


(a)

"\theta(0) = 3.04 \:\rm rad"

(b)

"\omega=\frac{d\theta}{dt} = 9.66t+ 8.97t^2"

"\omega(0) =0"

(c)

"\omega(7.43)= 9.66\times 7.43+ 8.97\times 7.43^2\\
=567\:\rm rad/s"

(d)

"\alpha=\frac{d\omega}{dt} = 9.66+ 17.9t"

"\alpha(1.12)= 9.66+ 17.9\times 1.12=29.8\:\rm rad/s^2"

(e) no